4.A student finds the %T of a 2.0 x 10-3M solution of Congo Red solution to be 35 % when measured in a 1.0-cm cuvet at a wavelength of 410 nm. What is the absorbance of this solution?
Answer #1

We will calculate the absorbance by using following equation,

Absorbance = log ( 100 / %T)

Here, log of 100 = 2.

Therefore,

Absorbance = 2 - log %T

= 2 - log 35 %

= 2 - 1.544

= 0.456

Therefore absorbance is 0.456.